OpenStudy (anonymous):

What percent of 360 is 270?

OpenStudy (lgbasallote):

so it's \[270 \div 360 \times 100\]

OpenStudy (anonymous):

\[ 270/360 × 100% = 0.75 × 100% = 75%\]

OpenStudy (anonymous):

\[= 0.75 × 100% = 75%\]

OpenStudy (anonymous):

=75